OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{−5|x + 1| = 10 }\) \(\normalsize\color{blue}{−5|x + 1|\color{red}{ \div (-5)} = 10 \color{red}{ \div (-5)} }\) \(\normalsize\color{blue}{|x + 1| = -2 }\) this gives that, \(\normalsize\color{blue}{x + 1 = -2 }\) or \(\normalsize\color{blue}{x + 1 = 2 }\)

OpenStudy (solomonzelman):

finish solving. Tell me if you need more help.

OpenStudy (anonymous):

@SolomonZelman Thanks can you also help me solve this Solve for x: -3|x + 7| = -12

OpenStudy (solomonzelman):

this one is solved similarly. Can you first divide both sides by `-3` ?

OpenStudy (anonymous):

@SolomonZelman Yes so it would be |x + 7|=4

OpenStudy (solomonzelman):

yes, and after that you know that ` │x+7│ `can be equal to 4, or it can be equal to -4. So you would write, \(\normalsize\color{blue}{ │x+7│=4~~~~~~~~\rm{or} ~~~~~~~~│x+7│=-4 }\)

OpenStudy (solomonzelman):

I mean \(\normalsize\color{blue}{ x+7=4~~~~~~~~\rm{or} ~~~~~~~~x+7=-4 }\)

OpenStudy (anonymous):

@SolomonZelman And x would be -3 or 3

OpenStudy (anonymous):

@SolomonZelman -3 or -11 i mean

OpenStudy (solomonzelman):

Yes, -11 and -3 :)

OpenStudy (anonymous):

@SolomonZelman Thanks :)

OpenStudy (solomonzelman):

Enjoy :)

OpenStudy (anonymous):

@SolomonZelman Hi again can you help me solve this one to Solve for x: |x − 2| + 10 = 12 Thanks